A sample of 16 ten-year-old girls had a mean weight of 71.5 and a standard deviation of 

12 pounds, respectively. Assuming normality, find the 90, 95, and 99 percent confidence 

intervals forthe mean of the population.


1
Expert's answer
2021-08-10T10:18:28-0400

"M=71.5 \\

\sigma=12 \\

n=16"

Two-sided confidence interval:

"CI = (M - \frac{Z_c \times \sigma}{\sqrt{n}}, M + \frac{Z_c \times \sigma}{\sqrt{n}})"

For 90 % confidence interval

"Z_c=1.645 \\

CI = (71.5 - \frac{1.645 \times 12}{\sqrt{16}}, 71.5 + \frac{1.645 \times 12}{\sqrt{16}}) \\

= (71.5 -4.935, 71.5+4.935) \\

= (66.565, 76.435)"

For 95 % confidence interval

"Z_c= 1.96 \\

CI = (71.5 - \frac{1.96 \times 12}{\sqrt{16}}, 71.5 + \frac{1.96 \times 12}{\sqrt{16}}) \\

= (71.5 -5.88, 71.5+5.88) \\

= (65.62, 77.38)"

For 99 % confidence interval

"Z_c= 2.576 \\

CI = (71.5 - \frac{2.576 \times 12}{\sqrt{16}}, 71.5 + \frac{2.576 \times 12}{\sqrt{16}}) \\

= (71.5 -7.728, 71.5+7.728) \\

= (63.772, 79.228)"